# Matrix-Free Two-to-Infinity and One-to-Two Norms Estimation

### v1

@@ 1 Introduction
In recent years, there has been growing interest in randomized linear algebra techniques (Halko, Martinsson, and Tropp 2011; Martinsson and Tropp 2020) for estimating matrix functions without explicit access to the matrix entries, see e.g. (Hutchinson 1989; Bekas, Kokiopoulou, and Saad 2007). This setting, known as matrix-free, assumes access to an oracle that computes matrix-vector products with a matrix $A$, and in some cases also with its transpose $A^{\top}$. The goal is to approximate important characteristics of $A$, for example, norms, trace or spectral density, using only these products. Such a framework is essential in modern machine learning, where matrices such as the Jacobian of deep neural networks are prohibitively large to form explicitly but allow efficient computation of matrix-vector products via automatic differentiation (autograd), see e.g. (Baydin et al. 2018).
@@
it suffices to concentrate on the estimation of the two-to-infinity norm. Compared to classical matrix norms such as the spectral or Frobenius norm, the two-to-infinity norm provides finer control over the row-wise structure of a matrix. Indeed, when dealing with tall matrices, that is, matrices $A\in\mathbb{R}^{d\times n}$ with $d\gg n$, it is natural that $\|A\|_{2\to\infty}$ does not scale with $d$, contrary to spectral and Frobenius norm of this matrix. In such cases, bounding the two-to-infinity norm ensures that the norm of each row is tightly controlled. This localized control is especially useful when studying theoretical properties of various algorithms (Cape, Tang, and Priebe 2019; Pensky 2024), see details in the related work section. In this paper, we focus on methodological and statistical aspects of estimating the two-to-infinity norm, rather than on its theoretical utility. Our main contributions are as follows:
@@
  • We apply suggested randomized estimators as regularizers in image classification and recommender systems problems. For image classification with deep neural networks, our method achieves better generalization performance compared to prior Jacobian regularization techniques (Hoffman, Roberts, and Yaida 2019; Roth, Kilcher, and Hofmann 2020). In the recommender systems domain, we consider UltraGCN-type architectures (Mao et al. 2021) and show that employing our regularizer increases the robustness of the algorithm to adversarial attacks, following the pipeline of (He et al. 2018).
@@ 2 Related Work
A classical problem in the matrix-free setting is the stochastic trace estimation (Hutchinson 1989), in which the trace of a matrix is approximated using a few matrix-vector products with random vectors. Several improvements on this approach have been developed, including variance-reduced methods such as Hutch++ (Meyer et al. 2021), which exploit low-rank structure to accelerate convergence, and dynamic algorithms (Dharangutte and Musco 2021; Woodruff, Zhang, and Zhang 2022), which adaptively allocate samples to achieve higher accuracy. Related work also addressed the problem of estimating the spectral norm of a matrix using structured estimators based on rank-one vectors, see (Bujanovic and Kressner 2021). Another closely related problem is the estimation of the matrix diagonal (Bekas, Kokiopoulou, and Saad 2007; Baston and Nakatsukasa 2022; Dharangutte and Musco 2023), for which recent works have provided algorithmic improvements and theoretical guarantees.
@@
The two-to-infinity norm is widely used as a theoretical tool to study statistical guarantees for algorithms in various areas of high-dimensional statistics and learning theory. Particular applications include bandit problems with specific reward structures (Jedra et al. 2024), singular subspace recovery (Cape, Tang, and Priebe 2019), and clustering (Pensky 2024).
@@
In machine learning algorithms, the two-to-infinity norm most often appears as part of the max-norm (see the notation section), an important tool for algorithm regularization. Its applications include, in particular, the matrix completion problem via matrix factorization (Srebro, Rennie, and Jaakkola 2004). In this setting, one can show that the max-norm serves as a surrogate for the rank of the approximator matrix. This idea was further developed in (Lee et al. 2010), where the authors introduce gradient-based algorithms, such as the projected gradient method and the proximal point method, for solving the regularized matrix factorization problem. Max-norm regularizers are also an important part of many online frameworks for such problems, often combined with additional sparsity-inducing constraints, as in (Shen, Xu, and Li 2014).
@@
Estimating matrix operator norms $\|A\|_{p\to q}$ induced by different combinations of $p$ and $q$ has recently attracted a lot of attention (Higham and Tisseur 2000; Bujanovic and Kressner 2021; Bresch et al. 2024; Naumov et al. 2025). A foundational contribution in this direction is the work (Boyd 1974), which proposes a general iterative algorithm for estimating $\|\cdot\|_{p\to q}$ norm for arbitrary $p$ and $q$, except for the cases of $\|\cdot\|_{1\to 2}$ and $\|\cdot\|_{2\to\infty}$ norms. Their approach is based on a generalization of the classical power method. Papers (Higham 1992; Roth, Kilcher, and Hofmann 2020) extend the methodology of (Boyd 1974) and apply it to more specialized norms, including the $\|\cdot\|_{1\to 2}$ and $\|\cdot\|_{2\to\infty}$ norm cases. While (Boyd 1974) proves theoretical convergence results for certain types of matrices, the authors of (Higham 1992; Roth, Kilcher, and Hofmann 2020) do not provide any theoretical guarantees for the $\|\cdot\|_{1\to 2}$ and $\|\cdot\|_{2\to\infty}$ cases. Furthermore, we provide an example below demonstrating that, when estimating the $\|\cdot\|_{2\to\infty}$ norm, this method diverges with positive probability even when applied to a simple diagonal matrix.
@@ 3 Two-to-Infinity Norm Estimation
Here, we briefly introduce the adaptive power method proposed by (Higham 1992; Roth, Kilcher, and Hofmann 2020). Full pseudocode is provided in the supplementary paper. The method begins with an initial random vector $X_{0}$ (typically drawn from a Gaussian distribution), followed by an iterative update process:
@@ 3.1 Hutchinson’s Diagonal Estimator
Unlike the iterative algorithms, such as the adaptive power method, we focus on a stochastic estimator of the two-to-infinity norm, which is accurate in the limit of large number of matrix-vector products. Our algorithms build upon a well-known technique for estimating the diagonal of a square matrix using only matrix-vector products. This technique is known as the Hutchinson diagonal estimator (Bekas, Kokiopoulou, and Saad 2007; Baston and Nakatsukasa 2022; Dharangutte and Musco 2023). In this section, we introduce the Hutchinson diagonal estimator and provide concentration inequalities that are useful for our subsequent analysis.
@@ 3.2 Our method
However, estimating the maximum value through the direct application of Hutchinson’s method introduces high variance, leading to a noisy approximation. To mitigate this, we can eliminate one source of randomness in the final estimate. Namely, let $D$ be the estimate of the diagonal of $AA^{\top}$. While the entries of $D$ are typically noisy, we can reduce variance by avoiding direct use of $\max_{i}D_{i}$. Instead, we first identify the (random) index of the maximum estimated value, $j=\arg\max_{i}D_{i}$, and then compute $\|A^{\top}e_{j}\|_{2}=\|A_{j}\|_{2}$. This approach significantly improves the quality of the estimate, which is supported by the ablation study carried out in Section 5.1. The outlined procedure is referred to as the TwINEst algorithm and is presented in Algorithm 1. Note that the algorithm steps given in lines $2$ to $4$ can be done either in parallel, or in a matrix form, if it is possible to store the matrix $X=[X^{1},\ldots,X^{m}]\in\mathbb{R}^{d\times m}$ and compute $X\odot AA^{\top}X$. This is the main practical advantage of our method over power-iteration based algorithms (Roth, Kilcher, and Hofmann 2020).
@@ Discussion.
The proof of Theorem 2 is provided in the supplementary paper. Importantly, the structure of TwINEst allows us not only to bound the probability of deviating from the true value, but also to bound the probability that our algorithm returns the exact value.
Remark 3 (TwINEst Sample Complexity).
@@
$m>\frac{2\log(2d/\delta)}{\varepsilon^{2}}\|AA^{\top}-\operatorname{diag}(AA^{\top})\|_{2\to\infty}^{2}$
to ensure $|T^{m}(A)-\|A\|_{2\to\infty}|<\varepsilon$ with probability at least $1-\delta$.
@@
Proof is completely analogous to the proof of Theorem 2. ∎
Depending on $\Delta$, either Remark 3 or Theorem 2 may be more useful. Results similar to those in Remark 3 have previously been obtained for the Hutchinson estimator (Roosta-Khorasani and Ascher 2015; Jiang et al. 2021), Hutch++ (Meyer et al. 2021), and several other methods. To the best of our knowledge, this is the first bound on the sample complexity of randomized estimation of the two-to-infinity norm.
@@ 4 Improved Algorithm
In this section, we present an improved version of our algorithm, which incorporates the variance reduction technique of the Hutch++ method (Meyer et al. 2021). Bearing similarity with the Hutch++ method, we call this modification TwINEst++. Rather than estimating the diagonal with the classical Hutchinson estimator, we adopt the unbiased Hutch++ diagonal modification (Han, Li, and Zhu 2024). The key idea of this technique is to approximate the dominant low‑rank part of the matrix using a few random vectors, explicitly compute its diagonal, and estimate the diagonal of the residual term using Hutchinson’s estimator. More precisely, let $S\in\mathbb{R}^{d\times r}$ consist of $r$ independent Rademacher vectors. Multiplying by $AA^{\top}$, we form the matrix $AA^{\top}S$, whose column span approximates the dominant eigenspace of $AA^{\top}$ with high probability. To approximate this eigenspace, we compute the thin $QR$-decomposition of the matrix $AA^{\top}S$ and further rely on the factor $Q\in\mathbb{R}^{d\times r}$. Using the orthogonal projector $P=QQ^{\top}$, we decompose $AA^{\top}$ into a low-rank approximation and a residual:
@@
$\operatorname{diag}(AA^{\top}P)=(A^{\top}(AQ)\odot Q)\cdot(1\ 1\ ...\ 1)^{\top},$
@@
This hybrid estimator typically achieves a significantly lower variance than the standard Hutchinson method, which in turn leads to better identification of the maximum row. A detailed description of the algorithm is provided in Algorithm 2. Below we provide the counterpart of Theorem 2 for the TwINEst++ estimator.
@@
7: Compute $\hat{D}=\frac{1}{m}\sum_{i=1}^{m}t_{i}\in\mathbb{R}^{d}$
@@
(a) Synthetic data with $\Delta=10^{-2}$.
(b) Synthetic data with $\Delta=10^{-1}$.
Theorem 4 shows that TwINEst++ achieves an improved oracle complexity compared to the original algorithm, particularly in challenging scenarios when $\Delta\to 0$, making identification of the correct row difficult. Specifically, the oracle complexity is reduced from $O(1/\Delta^{2})$ in the original TwINEst algorithm to $O(1/\Delta)$ in TwINEst++. However, due to an unknown constant in the TwINEst++ oracle complexity bound, the original TwINEst may perform better in practice for some matrices.
@@ 5 Experiments
In this section, we present an empirical evaluation of the proposed algorithms. We focus on comparing the algorithms on synthetic and real-world matrices (see Section 5.1 and Section 5.2), and applications of our methods to image classification and recommender systems tasks (see Section 5.3 and Section 5.4).
All experiments are conducted on a single NVIDIA Tesla V100 GPU with 32GB memory.
@@ 5.1 Synthetic Data
  • Adaptive Power Method. A modification of the power iteration method for estimating the two-to-infinity norm (see Algorithm 3 in the supplementary paper), based on (Higham 1992; Roth, Kilcher, and Hofmann 2020).
  • Rademacher Averaging. A version of the TwINEst method that does not compute the exact norm of the candidate row with the maximum norm. Formally, for a matrix $A\in\mathbb{R}^{d\times n}$ we output $\max_{i}D^{m}_{i}(AA^{\top})$ (see notations in Algorithm 1).
@@
For our synthetic experiments, we generate random Gaussian matrices $A\in\mathbb{R}^{5000\times 5000}$. Specifically, we fix a parameter $\Delta\in(0,1)$ and sample values $c_{3},\dots,c_{5000}\sim\mathcal{U}[0,1]$, setting $c_{1}=1+\Delta$, $c_{2}=1$. Each row of the matrix $A$ is normalized such that its squared $\ell_{2}$-norm is equal to $c_{i}$, ensuring a gap of magnitude $\Delta$ between the largest and second-largest squared row norms. The singular value distributions of these matrices are quite similar for different values of $\Delta$ (see Figure 4(a)). Therefore, the parameter $\Delta$ may have a significant impact on the convergence rate.
The results of synthetic experiments for different values of $\Delta$ are illustrated in Figure 1. As anticipated, TwINEst consistently outperforms Rademacher Averaging. The Adaptive Power Method from (Roth, Kilcher, and Hofmann 2020) fails to converge, as evidenced by its flat performance line. For a relatively large gap $\Delta=10^{-1}$, the TwINEst algorithm rapidly converges, achieving accurate results consistently within approximately 400 iterations.

### v2

@@ 1 Introduction
In recent years, there has been growing interest in randomized linear algebra techniques (Halko et al. 2011; Martinsson and Tropp 2020) for estimating matrix functions without explicit access to the matrix entries, see e.g. (Hutchinson 1989; Bekas et al. 2007). This setting, known as matrix-free, assumes access to an oracle that computes matrix-vector products with a matrix $A$, and in some cases also with its transpose $A^{\top}$. The goal is to approximate important characteristics of $A$, for example, norms, trace or spectral density, using only these products. Such a framework is essential in modern machine learning, where matrices such as the Jacobian of deep neural networks are prohibitively large to form explicitly but allow efficient computation of matrix-vector products via automatic differentiation (autograd), see e.g. (Baydin et al. 2018).
@@
it suffices to concentrate on the estimation of the two-to-infinity norm. Compared to classical matrix norms such as the spectral or Frobenius norm, the two-to-infinity norm provides finer control over the row-wise structure of a matrix. Indeed, when dealing with tall matrices, that is, matrices $A\in\mathbb{R}^{d\times n}$ with $d\gg n$, it is natural that $\|A\|_{2\to\infty}$ does not scale with $d$, contrary to spectral and Frobenius norm of this matrix. In such cases, bounding the two-to-infinity norm ensures that the norm of each row is tightly controlled. This localized control is especially useful when studying theoretical properties of various algorithms (Cape et al. 2019; Pensky 2024), see details in the related work section. In this paper, we focus on methodological and statistical aspects of estimating the two-to-infinity norm, rather than on its theoretical utility. Our main contributions are as follows:
@@
  • We apply suggested randomized estimators as regularizers in image classification and recommender systems problems. For image classification with deep neural networks, our method achieves better generalization performance compared to prior Jacobian regularization techniques (Hoffman et al. 2019; Roth et al. 2020). In the recommender systems domain, we consider UltraGCN-type architectures (Mao et al. 2021) and show that employing our regularizer increases the robustness of the algorithm to adversarial attacks, following the pipeline of (He et al. 2018).
@@ 2 Related Work
A classical problem in the matrix-free setting is the stochastic trace estimation (Hutchinson 1989), in which the trace of a matrix is approximated using a few matrix-vector products with random vectors. Several improvements on this approach have been developed, including variance-reduced methods such as Hutch++ (Meyer et al. 2021), which exploit low-rank structure to accelerate convergence, and dynamic algorithms (Dharangutte and Musco 2021; Woodruff et al. 2022), which adaptively allocate samples to achieve higher accuracy. Related work also addressed the problem of estimating the spectral norm of a matrix using structured estimators based on rank-one vectors, see (Bujanovic and Kressner 2021). Another closely related problem is the estimation of the matrix diagonal (Bekas et al. 2007; Baston and Nakatsukasa 2022; Dharangutte and Musco 2023), for which recent works have provided algorithmic improvements and theoretical guarantees.
@@
The two-to-infinity norm is widely used as a theoretical tool to study statistical guarantees for algorithms in various areas of high-dimensional statistics and learning theory. Particular applications include bandit problems with specific reward structures (Jedra et al. 2024), singular subspace recovery (Cape et al. 2019), and clustering (Pensky 2024).
@@
In machine learning algorithms, the two-to-infinity norm most often appears as part of the max-norm (see the notation section), an important tool for algorithm regularization. Its applications include, in particular, the matrix completion problem via matrix factorization (Srebro et al. 2004). In this setting, one can show that the max-norm serves as a surrogate for the rank of the approximator matrix. This idea was further developed in (Lee et al. 2010), where the authors introduce gradient-based algorithms, such as the projected gradient method and the proximal point method, for solving the regularized matrix factorization problem. Max-norm regularizers are also an important part of many online frameworks for such problems, often combined with additional sparsity-inducing constraints, as in (Shen et al. 2014).
@@
Estimating matrix operator norms $\|A\|_{p\to q}$ induced by different combinations of $p$ and $q$ has recently attracted a lot of attention (Higham and Tisseur 2000; Bujanovic and Kressner 2021; Bresch et al. 2024; Naumov et al. 2025). A foundational contribution in this direction is the work (Boyd 1974), which proposes a general iterative algorithm for estimating $\|\cdot\|_{p\to q}$ norm for arbitrary $p$ and $q$, except for the cases of $\|\cdot\|_{1\to 2}$ and $\|\cdot\|_{2\to\infty}$ norms. Their approach is based on a generalization of the classical power method. Papers (Higham 1992; Roth et al. 2020) extend the methodology of (Boyd 1974) and apply it to more specialized norms, including the $\|\cdot\|_{1\to 2}$ and $\|\cdot\|_{2\to\infty}$ norm cases. While (Boyd 1974) proves theoretical convergence results for certain types of matrices, the authors of (Higham 1992; Roth et al. 2020) do not provide any theoretical guarantees for the $\|\cdot\|_{1\to 2}$ and $\|\cdot\|_{2\to\infty}$ cases. Furthermore, we provide an example below demonstrating that, when estimating the $\|\cdot\|_{2\to\infty}$ norm, this method diverges with positive probability even when applied to a simple diagonal matrix.
@@ 3 Two-to-Infinity Norm Estimation
Here, we briefly introduce the adaptive power method proposed by (Higham 1992; Roth et al. 2020). Full pseudocode is provided in the supplementary paper. The method begins with an initial random vector $X_{0}$ (typically drawn from a Gaussian distribution), followed by an iterative update process:
@@ 3.1 Hutchinson’s Diagonal Estimator
Unlike the iterative algorithms, such as the adaptive power method, we focus on a stochastic estimator of the two-to-infinity norm, which is accurate in the limit of large number of matrix-vector products. Our algorithms build upon a well-known technique for estimating the diagonal of a square matrix using only matrix-vector products. This technique is known as the Hutchinson diagonal estimator (Bekas et al. 2007; Baston and Nakatsukasa 2022; Dharangutte and Musco 2023). In this section, we introduce the Hutchinson diagonal estimator and provide concentration inequalities that are useful for our subsequent analysis.
@@ 3.2 Our method
However, estimating the maximum value through the direct application of Hutchinson’s method introduces high variance, leading to a noisy approximation. To mitigate this, we can eliminate one source of randomness in the final estimate. Namely, let $D$ be the estimate of the diagonal of $AA^{\top}$. While the entries of $D$ are typically noisy, we can reduce variance by avoiding direct use of $\max_{i}D_{i}$. Instead, we first identify the (random) index of the maximum estimated value, $j=\arg\max_{\mathrm{i}}D_{\mathrm{i}}$, and then compute $\|A^{\top}e_{j}\|_{2}=\|A_{j}\|_{2}$. This approach significantly improves the quality of the estimate, which is supported by the ablation study carried out in Section 5.1. The outlined procedure is referred to as the TwINEst algorithm and is presented in Algorithm 1. Note that the algorithm steps given in lines $2$ to $4$ can be done either in parallel, or in a matrix form, if it is possible to store the matrix $X=[X^{1},\ldots,X^{m}]\in\mathbb{R}^{d\times m}$ and compute $X\odot AA^{\top}X$. This is the main practical advantage of our method over power-iteration based algorithms (Roth et al. 2020).
@@ Discussion.
The proof of Theorem 2 is provided in the supplementary paper. Importantly, the structure of TwINEst allows us not only to bound the probability of deviating from the true value, but also to bound the probability that our algorithm returns the exact value. Notably, when $\Delta=0$, identifying the maximum row norm becomes trivial, and the algorithm always returns the correct answer.
Remark 3 (TwINEst Deviation Bound).
@@
$m>\frac{8\log(2d/\delta)}{\varepsilon^{2}}\|AA^{\top}-\operatorname{diag}(AA^{\top})\|_{2\to\infty}^{2}$
to ensure $|(T^{m}(A))^{2}-\|A\|_{2\to\infty}^{2}|<\varepsilon$ with probability at least $1-\delta$.
@@
Proof is analogous to the proof of Theorem 2. ∎
Depending on $\Delta$, either Remark 3 or Theorem 2 may be more useful. Results similar to those in Remark 3 have previously been obtained for the Hutchinson estimator (Roosta-Khorasani and Ascher 2015; Jiang et al. 2021), Hutch++ (Meyer et al. 2021), and several other methods. To the best of our knowledge, Theorem 2 is the first bound on the oracle complexity of randomized estimation of the two-to-infinity norm.
@@ 4 Improved Algorithm
In this section, we present an improved version of our algorithm, which incorporates the variance reduction technique of the Hutch++ method (Meyer et al. 2021). Bearing similarity with the Hutch++ method, we call this modification TwINEst++. Rather than estimating the diagonal with the classical Hutchinson estimator, we adopt the unbiased Hutch++ diagonal modification (Han et al. 2024). The key idea of this technique is to approximate the dominant low‑rank part of the matrix using a few random vectors, explicitly compute its diagonal, and estimate the diagonal of the residual term using Hutchinson’s estimator. More precisely, let $S\in\mathbb{R}^{d\times r}$ consist of $r$ independent Rademacher vectors. Multiplying by $AA^{\top}$, we form the matrix $AA^{\top}S$, whose column span approximates the dominant eigenspace of $AA^{\top}$ with high probability. To approximate this eigenspace, we compute the thin $QR$-decomposition of the matrix $AA^{\top}S$ and further rely on the factor $Q\in\mathbb{R}^{d\times r}$. Using the orthogonal projector $P=QQ^{\top}$, we decompose $AA^{\top}$ into a low-rank approximation and a residual:
@@
$\operatorname{diag}(AA^{\top}P)=(A(A^{\top}Q)\odot Q)\cdot(1\ 1\ ...\ 1)^{\top},$
@@
This hybrid estimator typically achieves a significantly lower variance than the standard Hutchinson method, which in turn leads to better identification of the maximum row. We follow the Hutch++ strategy and set $r=m/3$, where $m$ is the whole budget for matrix-vector products (see ablation of different strategies for choosing $r$ in the supplement materials). A detailed description of the algorithm is provided in Algorithm 2. Below we provide the counterpart of Theorem 2 for the TwINEst++ estimator.
@@
7: Compute $\hat{D}=\frac{3}{m}\sum_{i=1}^{m/3}t_{i}\in\mathbb{R}^{d}$
@@
(a) Synthetic data with $\Delta=10^{-2}$.
(b) Synthetic data with $\Delta=10^{-1}$.
@@
Theorem 4 shows that TwINEst++ achieves an improved oracle complexity compared to the original algorithm, particularly in challenging scenarios when $\Delta\to 0$, making identification of the correct row difficult. Specifically, the oracle complexity is reduced from $O(1/\Delta^{2})$ in the original TwINEst algorithm to $O(1/\Delta)$ in TwINEst++. Moreover, we expect that TwINEst++ performs particularly well in low-rank settings, exploiting the low-rank structure of the matrices, as in recent methods (Meyer et al. 2021; Parkina and Rakhuba 2025). However, TwINEst’s oracle complexity depends on the norm of the matrix without its diagonal, while for TwINEst++ it depends on the norm of the whole matrix, so TwINEst may perform better for some matrices.
@@ 5 Experiments
In this section, we present an empirical evaluation of the proposed algorithms. We focus on comparing the algorithms on synthetic and real-world matrices (see Section 5.1 and Section 5.2), and applications of our methods to image classification and recommender systems tasks (see Section 5.3 and Section 5.4). All experiments are conducted on a single NVIDIA Tesla V100 GPU with 32GB memory.
@@ 5.1 Synthetic Data
  • Adaptive Power Method. A modification of the power iteration method for estimating the two-to-infinity norm (see Algorithm 3 in the supplementary paper), based on (Higham 1992; Roth et al. 2020).
  • Rademacher Averaging. A version of the TwINEst method that does not compute the exact norm of the candidate row with the maximum norm. Formally, for a matrix $A\in\mathbb{R}^{d\times n}$ we output $\sqrt{\max_{i}D^{m}_{i}(AA^{\top})}$ (see notations in Algorithm 1).
@@
For our synthetic experiments, we generate random Gaussian matrices $A\in\mathbb{R}^{5000\times 5000}$. Specifically, we fix a parameter $\Delta\in(0,1)$ and sample values $c_{3},\dots,c_{5000}\sim\mathcal{U}[0,1]$, setting $c_{1}=1+\Delta$, $c_{2}=1$. Each row of the matrix $A$ is normalized such that its squared $\ell_{2}$-norm is equal to $c_{i}$, ensuring a gap of magnitude $\Delta$ between the largest and second-largest squared row norms. The singular value distributions of these matrices are quite similar for different values of $\Delta$ (see Figure 4(a) in the supplementary paper). Therefore, the parameter $\Delta$ may have a significant impact on the convergence rate.
The results of synthetic experiments for different values of $\Delta$ are illustrated in Figure 1. As anticipated, TwINEst consistently outperforms Rademacher Averaging. The Adaptive Power Method fails to converge, as evidenced by its flat performance line. For a relatively large gap $\Delta=10^{-1}$, the TwINEst algorithm rapidly converges, achieving accurate results consistently within approximately 400 iterations.